OpenStudy (anonymous):

What is the y-intercept of 12x + 4y = 10

OpenStudy (anonymous):

5/2

OpenStudy (anonymous):

to do that, we have to make y the subject of the formula, so to do that, we must subtract 12x from both sides, and then divide every item or term by 4 so we have y=-3x+ (5/2) and y=mx+b b is the y intercept, so b = (5/2) or 2.5

OpenStudy (anonymous):

To get the y-intercept, put x=0 12(0) + 4y = 10 4y =10 y = 10/4 = 5/2 = 2.5